OpenStudy (anonymous):

4(10-5) ------- 3+2 4(10-5) ------- 3+2 @Mathematics

OpenStudy (saifoo.khan):

\[\frac{4(5)}{5}= \frac{20}{5}=4\]

OpenStudy (anonymous):

4(5)/5=20/5=4 is the correct answer